a small steel ball falls through a syrup at constant speed of 10 m/s. if steel ball is pulled upwards with a force equal to twice its effective weight, how fast will it move upwards

Grade:12

1 Answers

Vikas TU
7 years ago
From work energy theorem,
mgh = 0.5mv^2
v = root(2gh)
or
h = 100/2g => 5 meter.
Force with it has been pulled upwards = 2mg = ma => a = 2g m/s^2
0^2 = u^2 – 2*2g*5
u^2 = 200
u = 10root(2) m/s